OpenStudy (anonymous):

identify the slope intercept form of the line with slope -3 that passes through the point

OpenStudy (espex):

This isn't a complete question.

OpenStudy (anonymous):

identify the slope intercept form of the line with slope -3 that passes through (-1, 12). a. y=-3x+12 b. y=-1x+12 c. y=-3x+9 d. y=-1x+9

OpenStudy (espex):

Given those equations, which are in "slope" intercept form, what 'x' will give you your 'y' when you plug it in?

OpenStudy (anonymous):

found it, thanks!

OpenStudy (espex):

You're welcome.